The number of men and women among professors two academic areas (Science or Humanities) from a simple random sample of small colleges were counted, and the results are shown in the table below.

Science Humanities
Men 28 32
Women 12 28

Use the ?=0.01α=0.01 level of significance to determine whether or not the relationship between sex and academic area is statistically significant.

(a) ?2=χ2=

(b) Magic number =

(c) Is the relationship between sex and academic area statistically significant?

A. Yes
B. No
Answer #1

Solution: o Given that Science Total Humo nities men a - 28 b=32 60 atb Women d:28 40 ctd C-12 60 Total 100 EN ho brd a&C tev2 x² 2:7778 b Magical number: Degree of foeedom df = (0-1)(C-1) - (2-1)(2-1) =(1) (1) df = 1 2 xioul C - 6.635 Thug, magical